A monthly record of glucose level of patients was recorded and found to have a mean

of 98 mg/dL and a standard deviation of 10 mg/dL. If the data is normally distributed,

what percent of the patients

a) has blood glucose level higher than 100 mg/dL?

b) are above normal (glucose level ≥ 120 mg/dL)?

c) has blood glucose level between 120 and 125 mg/dL?


1
Expert's answer
2021-11-21T13:29:09-0500

a)


"P(X>100)=1-P(Z\leq\dfrac{100-98}{10})"

"=1-P(Z\leq0.2)\approx0.420740"


"42.0740\%"


b)


"P(X\geq120)=1-P(Z<\dfrac{120-98}{10})"

"=1-P(Z<2.2)\approx0.013903"

"1.3903\%"


c)



"P(120<X<125)=P(X<125)-P(X\leq120)"

"=P(Z<\dfrac{125-98}{10})-P(Z\leq\dfrac{120-98}{10})"

"=P(Z<2.7)-P(Z\leq2.2)"

"\approx0.9965330-0.9860966"

"\approx0.010436"

"1.0436\%"